OpenStudy (anonymous):

guys i need help on this exact DE equation: (x+2y)dx+(2x+3y+1)dy=0

OpenStudy (loser66):

@Hero

OpenStudy (mimi_x3):

so check it first \[M = x+ 2y => M_y =2\] \[N=2x+3y+1 => N_y=2\] Therefore, it's exact.

OpenStudy (anonymous):

can i know what is M and what is N?

OpenStudy (anonymous):

the answer that i got from my group is kinda different

OpenStudy (mimi_x3):

well your DE is \[(x+2y)dx +(2x+3y+1)d0\]

OpenStudy (mimi_x3):

I haven't gotten to the answer yet lol!

OpenStudy (anonymous):

xD ok sry

OpenStudy (mimi_x3):

\[f(x,y) = \int (x+2y)dx = \frac{x^2}{2}+2xy +h(y)\] then differentiate \(f(x,y)\) for \(h(y)\) (with respect to \(y\)) \[=> \frac{x^2}{2}+2xy+h(y) => 2x + g'(y)\] \[=> 2x+ g'(y) = 2x+ 3y +1\] so \[g'(y) = 3y+1\] integrating \[g'(y) = 3y+1\] with respect to \(y\) leads to \[g(y) = \int (3y+1)dy = \frac{3}{2}y^2+y\] so substituting \[\frac{x^2}{2}+2xy+ \frac{3}{2}y^2+y =c_1\]

OpenStudy (anonymous):

nice, thnx mimi for answering it, its exactly as the default answer and its logical

OpenStudy (mimi_x3):

pardon? default answer?

OpenStudy (anonymous):

i mean the answer which my group got, bah pardon my english, its not my main language

terenzreignz (terenzreignz):

You're actually pretty understandable :D

OpenStudy (anonymous):

mimi, are you a lecturer of some sort?

terenzreignz (terenzreignz):

No, she's just awesome :D

OpenStudy (anonymous):

:) glad there are people like her

terenzreignz (terenzreignz):

Welcome to Openstudy, @5alood Stick around, maybe there's somebody with a question you can help with ^.^

OpenStudy (anonymous):

thank you, this is really helpful, and i will do my best to help :)